The reaction of 5.75 g of carbon with excess O2 yields 11.4 g of CO2. What is the percent yield of this reaction?
Answer #1

C + O2 ----------> CO2

moles of C = mass/molar mass = 5.75/12 = 0.4792 moles

From reaction,

1 mole C produces 1 mole CO2

So, 0.4792 mole C will produce 0.4792 moles of CO2

So, mass of CO2 = Moles*molar mass = 0.4792*44 = 21.0848 grams.

% yield of CO2 = 11.4*100/21.0848 = 54.0674 % .... Answer